The focus is at (-1,7), the length from the focus to the vertex is 2 units, and the function has a minimum.
Write the equation of the parabola.

Answer by lwsshak3(11628)

The focus is at (-1,7), the length from the focus to the vertex is 2 units, and the function has a minimum. Write the equation of the parabola.
***
Parabola opens upwards: (curve has a minimum)
Basic equation of parabola: (x-h)^2=4p(y-k)
p=2 (distance from vertex to focus)
4p=8
h=-1(same as x-coordinate(-1) of focus)
k=5(2 units below y-coordinate(7) of focus)
Equation:
(x+1)^2=8(y-5)